OpenStudy (katielong):

show that integating lnx/x^3 with limits of e and 1, equals (e^2 -3)/4e^2

OpenStudy (katielong):

\[\int\limits_{1}^{e}\frac{ lnx }{ x^3 } = \frac{ e^2 -3 }{ 4e^2 }\]

OpenStudy (katielong):

to begin with, i've integrated by parts, using: u=lnx du/dx = 1/x v=\[-\frac{ 1 }{ 2 } x ^{-2}\] dv/dx= \[x^-3\]

OpenStudy (irishboy123):

that's good do far

OpenStudy (katielong):

i've got: \[\int\limits_{1}^{e}f(x) dx= -\frac{ 1 }{ 2 }x ^{-2} lnx +\frac{ 1 }{ 2 } \int\limits_{1}^{e} x ^{-3} dx\]

OpenStudy (irishboy123):

yup

OpenStudy (irishboy123):

but first bit is within interval too -- pedantic

OpenStudy (katielong):

yeah sorry ahaha, so then i get...

OpenStudy (katielong):

\[\left[ \frac{ 1 }{ 4 } e ^{-2} (2lnx-1)\right]\ \lim_{e \rightarrow 1}\]

OpenStudy (irishboy123):

mmm you mean x^-2 AND check your signs....

OpenStudy (katielong):

yes sorry, and which sign?

OpenStudy (katielong):

.. and how on earth do i get it into the form stated in the question?

OpenStudy (irishboy123):

get the solution right first. +1/2 ∫ x^-3 dx = MINUS 1/4 x^-2 both terms in solution have a minus sign

OpenStudy (irishboy123):

when you have the solution done correctly, knowing that ln e = 1 & ln 1 = 0 gives you the answer but get the solution all lined up before doing that...

OpenStudy (katielong):

i cant get it into that form, im stuck i've got \[(-\frac{ 1 }{ 4e ^{2} }(3)) - (-\frac{ 1 }{ 4})\]

OpenStudy (katielong):

=1/4 - 3/4e^2

OpenStudy (irishboy123):

the final integral is \[-\frac{lnx}{2x^{2}} - \frac{1}{4x^{2}}\] it simplifies to \[-\frac{1}{4x^2}(2lnx + 1)\]

OpenStudy (irishboy123):

you are almost done =1/4 - 3/4e^2 =1/4 * (e^2 / e^2) - 3/4e^2 = e^2/ 4 e^2 - 3 / 4 e^2 = (e^2 - 3) / 4 e^2

OpenStudy (katielong):

ohhhh now i see it! you have to make the denomenators the same to add them, ahaha thank you very much!